The length of a rectangle is 4 inches less than 3 times the width and the perimeter is 32. Find the dimensions of the rectangle.
Answer by ptaylor(2198)

Perimeter(P)=2*Length(L) +2*Width(W) or P=2L+2W
Let W=width
Then Length=3W-4
Soooo
32=2(3W-4)+2W simplify
32=6W-8+2W collect like terms and add 8 to each side
32+8=8W
8W=40
W=5
Length=3*5-4=11
CK
2*11+2*5=32
Hope this helps--ptaylor